OpenStudy (anonymous):

If g(x) = x2 + 3, find g(4). (2 points) 11 19 16 8

OpenStudy (ronlover101):

Substitute the values. 3. g(x) = x^2 + 3, find g(4) So, 4 = 4^2 + 3. Can you simplify from there? 4. If h(x) = 2x + 1, find the value of x if h(x) = 11 So, 11 = 2x + 1 and solve for x.

OpenStudy (ronlover101):

Thanks!